OpenStudy (kmullis6):

An equation is shown below: 4(x - 3) - 5(x + 1) = 3 Which statement shows a correct next step in solving the equation? The equation can become 4x - 3 - 5x + 1 = 3 by applying the associative property of multiplication. The equation can become 4x - 3 - 5x + 1 = 3 by applying the distributive property. The equation can become 4x - 12 - 5x - 5 = 3 by applying the distributive property. The equation can become 4x - 12 - 5x - 5 = 3 by applying the associative property of multiplication. ---------- What is distributive and associative properties?

OpenStudy (texaschic101):

The equation can become 4x - 12 - 5x - 5 by using the distributive property. The distributive property : a (b + c) = ab + ac...basically, you just multiply everything in the parenthesis by the number outside of the parenthesis. The associative property : (a x b) x c = a x (b x c)....its basically saying that it does not matter how you group the numbers, the answer will be the same. any questions ?

OpenStudy (kmullis6):

No, that cleared it up for me. Thanks so much! @texaschic101

OpenStudy (texaschic101):

no problem :)

OpenStudy (anonymous):

The equation can become 4x - 12 - 5x - 5 = 3 by applying the distributive property.

OpenStudy (anonymous):

3 + 4x = 5x + -4 3 + 4x = -4 + 5x Solving for variable 'x'.